OpenStudy (anonymous):

A can do a work in 12 days ;B in 6 days and c in 3 days.A and B start working together and after a day,c joins them .the total no of days required to complete the work

OpenStudy (anonymous):

speak in english i cant understand

OpenStudy (anonymous):

wat u wrote before@uri

OpenStudy (debbieg):

The amount of the job completed by each is the reciprocal of how long it takes each to do the job. So if R=rate, we can say Ra=1/12 (a does 1/12 of the job in a day) Rb=1/6 (b does 1/6 of the job in a day) Rc=1/3 (c does 1/3 of the job in a day) So let x=# of days a & b work. Then an expression for the # of days that c works is....???

OpenStudy (debbieg):

Now the amount of the job done by each is their rate x their number of days worked. You need all of that to add up to 1 (which is "1 whole job" completed.) Then solve for x! :)

OpenStudy (uri):

Debbie..look closely..carefully.

OpenStudy (debbieg):

What @uri ??

OpenStudy (anonymous):

1/4=a+b; a+b+c=7/12;

OpenStudy (anonymous):

a+b+c=7/12

OpenStudy (anonymous):

1and 2/7

OpenStudy (anonymous):

plse help me

OpenStudy (anonymous):

A: has a capacity of 1/12 works per day B: has a capacity of 1/6 works per day C: has a capacity of 1/3 works per day In order to complete one work, A and B work n days and C n-1 days, then you have: n/12+n/6+(n-1)/3=1 work, then: n/12+2n/12+(4n-4)/12=1---->n+2n+4n=12+4---->7n=16---->n=16/7 days

OpenStudy (anonymous):

Test the solution: 16/7*(1/12+1/6)+(16/7-1)*1/3=1

OpenStudy (debbieg):

Sorry @jimra, I missed the notifications earlier. Carlos has done the whole problem for you above, just as I was trying to describe to you how to do it, so make sure you understand everything he wrote. :)

OpenStudy (anonymous):

Sorry for that @DebbieG I did not realize you had already outlined the solution. The medal is for you